OpenStudy (anonymous):

Solve using quadratic formula 6p^2=-25p-4

OpenStudy (anonymous):

\[6p ^{2}+25p+4=0\]

OpenStudy (anonymous):

p=[-25(+/_)sqrt(625-96)]/2

OpenStudy (anonymous):

p=[-25(+/-)sqrt(529)]/12 ohi missed 12 in the previous post

OpenStudy (anonymous):

p=[-25(+/-)23]/12 p=-25+23/12=-2/12=-1/6 or p=-25-23/12=-48/12=-4